OpenStudy (anonymous):

calculate plane equation (1,0,0), (0,1,0),(0,0,1)

OpenStudy (amistre64):

yay.... given three point we can create 2 vectors and cross product for the noral right?

OpenStudy (anonymous):

yes. but please do it. i'm being too lazy =P

OpenStudy (amistre64):

LOL.... happy to, i need the practice ;)

OpenStudy (amistre64):

lets choose a starting point; (1,0,0) looks fine; and create a vector from that point another and then the other

OpenStudy (anonymous):

cheers!

OpenStudy (amistre64):

(1,0,0) (0,1,0) ------- <1,1,0> = v1 (1,0,0) (0,0,1) ------ <1,0,1> = v2 right?

OpenStudy (amistre64):

we cross these 2 vectors to get a normal vector to the plane: < i, j, k> <1,1,0> <1,0,1> ---------- +(1-0) | - (1-0) | = <1,1,-1> for the normal +(0-1) |

OpenStudy (anonymous):

just the pane eq...

OpenStudy (amistre64):

now attach the normal to a point... once you fix my typo lol <1,-1,-1> = n ; and P(1,0,0) 1(x-1) -1(y+0) -1(z+0) = 0

OpenStudy (amistre64):

x-1 -y -z = 0 x-y-z-1 = 0 should satisfy it

OpenStudy (amistre64):

i wonder if i did my point addition right...

OpenStudy (anonymous):

thanks alot1 =)

OpenStudy (amistre64):

was it correct?

OpenStudy (anonymous):

i dunno =P its part of a long question.

OpenStudy (amistre64):

well, let me dbl check me thoughts and verify my results :)

OpenStudy (anonymous):

yes plz =)

OpenStudy (anonymous):

it correct! thumbs up!

OpenStudy (amistre64):

yay!! ..... i was like 84% sure :)

OpenStudy (anonymous):

x+y+z-1=0 the normal vector shd be 1,1,1. Use intuition. If you want to use the cross product to find the normal vector, it shd be: (1,-1,0)×(0,1,-1)=(1,1,1). Bear in mind that you should subtract one vector from another, NOT add them together. The next step should be verifying your answer, in this case, try sub. (1,0,0), (0,1,0), (0,0,1) into your equation to see if it makes sense.

OpenStudy (amistre64):

thats what my doubts were :) I saw that in the back of my head ..

OpenStudy (amistre64):

but i couldnt seem to get the crossproduct of the vectors; even with subtraction; to conform...

OpenStudy (amistre64):

(1,0,0) (1,0,0) -(0,1,0) -(0,0,1) ------- ------ <1,-1,0> <1,0,-1> <1,-1,0> <1,0,-1> <1-0,-(-1-0),0-1> = <1,1,-1>

OpenStudy (amistre64):

1(x-1) +1(y+0)-1(z+0) = 0 x + y -z -1 = 0 ....

OpenStudy (amistre64):

(0,0,1) = -2 then

OpenStudy (anonymous):

So the true answer is x+y+z=1 for your k component in the post 4 min ago, it should be (0)(1)-(-1)(1)=1

OpenStudy (amistre64):

.... i forgot how to multiply lol....

OpenStudy (anonymous):

^.^ Are you from USA?

OpenStudy (amistre64):

I am

OpenStudy (amistre64):

i kept breezing over that k part and needed a 2nd set of eyes to overcome my inherent stupidity :) thanx..

OpenStudy (anonymous):

"add oil" Guess from where I come from.

OpenStudy (amistre64):

venezula?

OpenStudy (amistre64):

cape fear?

OpenStudy (amistre64):

madagascar... gotta be madagascar :)

OpenStudy (anonymous):

no =] it should not be hard to guess if you google "add oil"

OpenStudy (amistre64):

its cantonese; soo... you from canton?

OpenStudy (anonymous):

Hong kong actually

OpenStudy (anonymous):

so add oil

OpenStudy (amistre64):

thnx :) you too